OpenStudy (anonymous):

A circle is described by the equation (x−12)2+(y−32)2=49. What are the cordinates for the center of the circle and the length of the radius?

OpenStudy (anx):

For a circle equation the center is always noted in the equation (x-12)^2 + (y-32)^ = 49 The center is placed in this format (x - h )( y - k) = r^2 so for (x-12)^2 and (y -32) the center is at (12, 32)